### v1

# Hierarchical Vision-Language Planning for Multi-Step Humanoid Manipulation

Abstract
Enabling humanoid robots to reliably execute complex multi-step manipulation tasks is crucial for their effective deployment in industrial and household environments. This paper presents a hierarchical planning and control framework designed to achieve reliable multi-step humanoid manipulation. The proposed system comprises three layers: (1) a low-level RL-based controller responsible for tracking whole-body motion targets; (2) a mid-level set of skill policies trained via imitation learning that produce motion targets for different steps of a task; and (3) a high-level vision-language planning module that determines which skills should be executed and also monitors their completion in real-time using pretrained vision-language models (VLMs). Experimental validation is performed on a Unitree G1 humanoid robot executing a non-prehensile pick-and-place task. Over 40 real-world trials, the hierarchical system achieved a 72.5% success rate in completing the full manipulation sequence. These experiments confirm the feasibility of the proposed hierarchical system, highlighting the benefits of VLM-based skill planning and monitoring for multi-step manipulation scenarios. See https://vlp-humanoid.github.io/ for video demonstrations of the policy rollout.
I Introduction
Humanoid robots offer a promising path towards general-purpose autonomy, integrated seamlessly into human environments. The humanoid morphology, while advantageous for tool use and ergonomic compatibility, presents challenging control and planning problems [12]. Recent reinforcement-learning (RL) and imitation-learning (IL) breakthroughs have produced striking single-skill demonstrations such as walking, dancing, or boxing [6, 11, 15, 14, 16]. Yet, practical applications require that the humanoid carry out multi-step tasks, i.e., interacting with multiple objects and regions over time. The sequencing and integration of these learned skills remains challenging, due to the high-dimensional nature of humanoid control, the requirement for continuous real-time visual feedback, and the complexity of determining precise transitions between consecutive skills.
To bridge the gap between isolated skill execution and long-horizon autonomy, robots must combine low-level motor dexterity with higher-level reasoning and task monitoring. Vision-Language Models (VLMs) present a unique opportunity to integrate planning and monitoring with robust learned control policies, enabling embodied humanoid agents to execute extended, multi-step manipulation tasks in unstructured, human environments [8, 29, 28].
In this paper, we present a hierarchical framework for humanoid manipulation planning and control, spanning three layers. The lowest level employs a whole-body tracking controller trained via Proximal Policy Optimization (PPO) [26], which tracks motion targets provided by a mid-level imitation learning (IL) policy. This mid-level policy is trained on human teleoperation data, where human poses are retargeted onto the robot’s morphology. Each IL-based skill policy maps egocentric visual observations and proprioceptive states to full-body motion commands, enabling the execution of complex loco-manipulation tasks. At the highest level, we propose a hierarchical planning module consisting of two VLMs: a low-frequency VLM Planner is used to generate sequences of IL skills from visual and textual task inputs, and a higher-frequency VLM-based skill-execution monitor that continuously verifies skill completion, orchestrating transitions between skills.
We empirically validate our proposed hierarchical system on the 29-DoF Unitree G1 humanoid robot, performing a representative multi-step pick-and-place task. In real-world testing, the integrated system successfully completed the full manipulation sequence in 72.5% of 40 trials. Our hierarchical approach demonstrates effective coordination of manipulation skills and visual feedback, highlighting the benefit of structured skill decomposition and VLM-based execution monitoring for reliable, multi-step humanoid autonomy.
The primary contributions of our work are: (1) a novel hierarchical vision-language planning and monitoring framework that dynamically sequences and verifies humanoid manipulation skills; and (2) an integrated autonomous system validated on a real humanoid robot performing complex multi-step tasks in realistic environments, demonstrating practical feasibility and effectiveness.
II Related Work
II-A Hierarchical Control in Humanoids
Recent approaches to humanoid control commonly split the control task into two separate layers: a low-level reinforcement learning (RL) controller that tracks reference motion targets, and a higher-level module that generates these motion targets. ExBody [6] first showed that a PPO tracker trained in Isaac Gym [23] can reproduce upper-body MoCap clips on the Unitree H1 while leaving the legs free for balance. ExBody2 [18] extends the idea to whole-body loco-manipulation via teacher-student distillation and domain randomization, achieving high-fidelity whole-body tracking on both H1 and G1 platforms. HumanPlus [11] replaces the multilayer perceptron (MLP) backbone of the low-level policy with a transformer backbone. Additionally, it expands the training data to use the AMASS dataset [22], significantly increasing the diversity of available human motion examples.
While low-level policies reliably track motion targets, generating these references is equally critical for accomplishing manipulation tasks. H2O (Human-to-Humanoid) [17] utilized single-camera human pose estimation to enable zero-shot teleoperation. OmniH2O [14] extended H2O’s concept, offering a unified interface for teleoperation using various modalities, including RGB cameras, VR, motion capture suits, and exoskeleton systems. OmniH2O further incorporated diffusion-based imitation learning [7], enabling the autonomous acquisition of manipulation skills directly from human demonstrations. Similarly, HumanPlus [11] adopted this hierarchical structure, leveraging human-teleoperated demonstrations recorded along with proprioceptive and egocentric visual data. These demonstrations were used to train the Humanoid Imitation Transformer (HIT)—a transformer-based imitation learning policy inspired by the Action Chunking Transformer (ACT) [31]. Using only 20–40 demonstrations per skill, HumanPlus reported 60–100
Across these studies, a common hierarchical structure appears, featuring a robust RL-based low-level policy for tracking motion references and a high-level module generating these references through teleoperation or imitation learning. Although this approach has shown success in single-skill tasks (e.g., dancing, lifting, and boxing), existing systems generally require human intervention when transitioning between skills, thus limiting their ability to operate autonomously over extended task sequences.
II-B VLMs for Long-Horizon Manipulation
Executing multi-step tasks demands semantic understanding, robust memory, and resilience to errors. Recent approaches leveraging large vision-language-action (VLA) models, such as RT-2 [4], OpenVLA [19], and $\pi_{0.5}$ [3], demonstrate promise in end-to-end planning but typically require large-scale paired datasets and lack interpretability. To address these limitations, hierarchical approaches integrate intermediate semantic representations, bridging abstract instructions and concrete actions (e.g., RT-H [2], NaVILA [5]). Another direction employs explicit hierarchies, where high-level vision-language models predict coarse geometric plans refined by lower-level controllers, effectively separating semantic reasoning from robot dynamics [21, 27]. Additionally, methods like KALM [10], RoboPoint [30], and ManipGen [9] use VLMs to predict spatial affordances or parameterized skill libraries, facilitating efficient zero-shot generalization.
Despite the advances in humanoid control, current humanoid systems lack integrated high-level modules capable of autonomously selecting and sequencing skills, as well as verifying successful execution. Addressing this gap, our work proposes an additional third hierarchical layer, combining a VLM-based planner and skill monitor on top of the established two-layer humanoid control stacks, enabling fully autonomous execution of extended loco-manipulation tasks.
III Method
Figure 2: Overview of the proposed hierarchical framework for autonomous multi-step humanoid manipulation. The system consists of: (1) a high-level planning module using vision-language models (VLMs) to generate task-specific skill sequences and monitor their execution via visual feedback; (2) a mid-level motion-generation module composed of imitation-learned action expert policies, mapping egocentric visual inputs and proprioceptive robot states into target joint configurations; and (3) a low-level reinforcement-learning-based tracking policy generating feasible joint-angle trajectories, executed by a high-frequency PD controller.
We propose a three-layer hierarchical control and planning framework for multi-step humanoid manipulation, illustrated in Fig. 2. The system consists of: (1) a low-level whole-body tracking controller trained via reinforcement learning (Sec. III-A), (2) a mid-level motion-generation module composed of multiple imitation-learned action expert policies, each trained from teleoperated demonstrations (Sec. III-B);, and (3) a high-level vision-language-based planning and execution monitoring module (Sec. III-C). The following sections detail each component.
III-A Low-Level Tracking Policy
Our low-level module is a reinforcement learning-based tracking controller designed to track generated motion goals. Building upon the ExBody architecture [6], we extended the implementation to the full 29-DoF Unitree G1 humanoid robot, incorporating whole-body tracking.
III-A1 Problem Formulation
The tracking policy, denoted as
$\pi_{T}:G\times S\rightarrow A,$
[…]
III-A2 Policy Architecture and Training
We employ a standard actor-critic architecture trained via Proximal Policy Optimization (PPO) [26]. Both the actor and critic networks are implemented as three-layer MLPs (256 hidden units, ReLU activations). Training occurs within the Isaac Gym simulator using 4096 parallel environments. To ensure robustness in real-world conditions, we extensively randomize environment parameters, including gravity, friction, mass distribution, motor strength, external perturbations, and terrain irregularities.
The reward function combines tracking accuracy terms, adapted from ExBody [6] and ExBody2 [18], along with additional stability-promoting regularization terms specifically tuned for the G1 robot. Detailed reward components are provided in Sec. -A.
III-A3 Training Data
Training data was sourced from a curated subset of the AMASS dataset [22], selected to cover the robot’s operational range. The data was supplemented by additional teleoperation demonstrations that we collected, tailored to the specific downstream skills and tasks outlined in the project. This combination ensures both diversity and task-relevance of the training data. Quantitative and qualitative evaluations of the tracking performance of the low-level policy outlined in Sec. -A.
III-B Mid-Level Imitation Learned Skills
Building upon the low-level tracking policy’s capability to execute motion goals precisely, the mid-level module generates higher-level skill behaviors autonomously. Specifically, it translates sensory inputs into coherent joint-motion targets that serve as reference trajectories for the tracking policy. This module integrates two complementary capabilities: an RGB-based teleoperation pipeline for demonstration collection, and imitation-learned (IL) expert policies trained from these demonstrations to execute individual skills autonomously.
III-B1 Teleoperation via Human Pose Retargeting
To facilitate the collection of demonstration data, we developed an RGB-based teleoperation pipeline allowing human operators to control the humanoid robot directly. This system consists of three primary stages: human pose estimation from RGB images, retargeting of human poses onto the robot, and generation of robot-specific keypoints.
For human pose estimation, we employ HybrIK (Hybrid Inverse Kinematics) [20], a hybrid analytical-neural inverse kinematics model. HybrIK initially predicts 3D joint positions from single-view RGB inputs using a convolutional neural network (CNN). Subsequently, these predictions are refined through a physics-informed inverse kinematics optimization step, yielding temporally smooth and anatomically plausible human joint configurations. This pipeline eliminates the need for specialized motion-capture equipment, requiring only a standard RGB camera setup.
Due to morphological differences between human operators and the Unitree G1 robot—including limb proportions, joint limits, and degrees of freedom—a dedicated retargeting procedure translates the human pose estimates into robot-compatible joint configurations. This retargeting procedure is outlined in Sec. -A.
As a final step, the low-level tracking policy expects as input both joint-angle references and corresponding 3D keypoints representing the center-of-mass (CoM) positions of selected robot body links. To generate these keypoints, we apply forward kinematics in simulation: we first set the robot to the retargeted joint configuration, and then compute the spatial coordinates of the predefined body positions. The resulting joint angles and keypoints form the complete motion reference that can be directly executed by the tracking controller.
III-B2 Imitation Learning Module
Building upon the established teleoperation pipeline, we next develop the imitation learning (IL) module, the mid-level component responsible for generating autonomous motion targets from onboard sensory inputs.
While the low-level tracking policy executes externally provided motion references, the IL module produces these references directly from sensory inputs, enabling robot autonomy. Formally, the IL policy is defined as a function
$\pi_{IL}:\mathcal{X}\rightarrow\mathcal{A},$
where the observation state at each timestep $t$ is
$x_{t}=(\mathbf{q}_{t},\,I_{t}^{(1)},\,I_{t}^{(2)}),$
comprising proprioceptive joint angles $\mathbf{q}_{t}$ and binocular egocentric RGB images $I_{t}^{(1)}$ and $I_{t}^{(2)}$. The action space $\mathcal{A}$ consists of future joint configurations $\mathbf{q}$, directly interfacing with the low-level tracking policy.
For the IL policies, we adopt the Humanoid Imitation Transformer (HIT) [11], designed specifically for high-DoF humanoids with binocular vision. It is based on the Action Chunking Transformer (ACT) [31] model, which have demonstrated improved robotic control by predicting multi-step future actions in one forward pass. At inference, HIT outputs action chunks of 50 joint-angle targets at 25 Hz. During training, an auxiliary $L_{2}$ loss between predicted and actual future visual embeddings improves visual grounding and generalization. The predicted joint angles are subsequently transformed into corresponding 3D keypoints via forward kinematics, matching the inputs required by the low-level tracking controller.
Instead of using the robot’s onboard Intel RealSense camera, we employ two externally mounted ELP high-speed RGB cameras with wide-angle lenses. This binocular configuration provides consistent frame rates and explicit depth information, enhancing spatial reasoning capabilities essential for manipulation tasks.
Demonstration data for IL training is collected through teleoperation, forming a dataset
$\mathcal{D}=\{(x_{1},a_{1}),(x_{2},a_{2}),\dots,(x_{N},a_{N})\},$
containing synchronized state-action pairs recorded at 25 Hz. We collect data for the representative task rearrangement involving picking a bag from a table and and placing it onto another. Only successful sequences (30 in total) are retained, segmented into separate skill-specific datasets, i.e. picking and placing, for training two policies. IL policy training employs the demonstration data with hyperparameters specified in Table V. After training, we evaluate independently every skill policy, evaluation over 30 autonomous trials showed success rates of $90\%$ for picking and $83.3\%$ for placing.
III-C Multi-Step Skill Planning and Execution Monitoring
While imitation-learned skills enable reliable execution of short-term actions, real-world humanoid tasks frequently require combining multiple skills into longer sequences to achieve complex goals. We address this by introducing a high-level planning and execution monitoring module, capable of dynamically selecting and verifying the execution of skill sequences. Our approach leverages pretrained Vision-Language Models (VLMs) embedded within a closed-loop planner-monitor architecture.
To clearly distinguish conceptual terminology, we define:
  • A skill as an individual, short-duration manipulation capability (e.g., picking or placing an object).
  • A task as a higher-level objective involving at least two skills executed sequentially that change the state of objects in the world.
Our hierarchical system thus plans at the task-level while monitoring and executing at the skill-level.
III-C1 Problem Formulation
Formally, a task is defined by a natural-language goal $g$ (e.g., “Pick up the bag and place it on the table”) and an initial visual observation $o_{0}$. The goal is to autonomously generate and reliably execute a sequence of parameterized skills $\sigma=[\pi^{(1)},\pi^{(2)},\dots,\pi^{(N)}]$ to satisfy the conditions described by $g$.
III-C2 System Architecture
Our high-level planning and monitoring module consists of two complementary components:
  1. (a)
    VLM Planner ($\mathcal{P}$): A GPT-4o model [24] generates structured, interpretable skill sequences from visual and textual task inputs.
  2. (b)
    VLM Skill Monitor ($\mathcal{M}$): A lightweight Gemini-2.0-Flash-Lite model [28] continuously verifies the completion of each executed skill at approximately 1 Hz.
Together, these two components form an iterative planning-monitoring loop, ensuring coherent multi-step task execution.
III-C3 Skill Library and PDDL-like Representation
We represent each skill available to the robot using structured, human-readable descriptions that include the ‘preconditions’ and ‘effects’ fields of an operator definition in Planning Domain Definition Language (PDDL) [1]. Each skill description includes a one-sentence summary, explicit natural-language preconditions, effects, and example verification queries, as illustrated for the Pick skill in Figure 3. Unlike rigid symbolic representations, this PDDL-like form balances logical rigor and flexibility, enabling the planner to perform common-sense reasoning without exhaustive state enumeration.
Pick:
Description: Pick up an object and hold it up
Preconditions: [
hand is empty’,
object is graspable’,
object is on a surface
Effects: [
robot is holding object up’,
object is no longer on the surface
Examples: Has the robot successfully picked up and is now holding the bag?’
Figure 3: Example structured skill description provided to the VLM Planner and Skill Monitor.
III-C4 VLM Planner
Given an initial image $o_{0}$, task instruction $g$, and the structured skill library, the VLM Planner $\mathcal{P}$ outputs a grounded sequence of executable skills $\sigma$. GPT-4o generates the skill sequence by visually grounding skill preconditions and effects, translating them into binary visual-question-answering queries for logical coherence. For instance, the planner implicitly understands causal ordering, such as the necessity of picking up an object before placing it elsewhere, from the provided natural-language annotations, without explicit symbolic planning constraints.
III-C5 VLM Skill Monitor
The VLM Skill Monitor $\mathcal{M}$ continuously evaluates the execution state through real-time visual feedback. Operating at approximately 1 Hz, it assesses short video snippets (10–15 frames extracted from 1.5-second segments) captured during skill execution. For each skill, the monitor uses the provided natural-language verification queries to output a binary decision: either completed or in progress. This approach ensures timely and reliable skill transitions while maintaining real-time responsiveness.
A walkthrough example of the functioning of this module is presented in Appendix -C while Appendix -D shows example VLM prompts.
III-D Full Hierarchical System
The complete integrated system implements the hierarchical pipeline presented in Figure 2. The pipeline comprises: (1) a high-level VLM Planner (GPT-4o) that generates a sequence of skills from initial visual inputs and natural-language task instructions; (2) a real-time VLM Skill Monitor (Gemini-2.0-Flash-Lite) that verifies skill execution status based on visual feedback; (3) a Skill Selector that dynamically chooses the current skill to execute based on the planned sequence and monitor feedback; (4) a mid-level Imitation Learning Module that activates the corresponding imitation-learned policy according to the selected skill, producing joint-motion targets from egocentric images and proprioceptive data; and (5) a low-level Tracking Policy combined with a high-frequency PD Controller, converting these joint targets into executable motor torque commands.
A key advantage of this hierarchical design is its inherent modularity: individual IL skills are trained independently and added directly to the skill library, enabling effortless expansion of the robot’s behavioral repertoire. Newly acquired skills become immediately available for integration by the high-level planning module, facilitating seamless composition into increasingly complex, multi-step tasks without extensive system reconfiguration.
IV Experiments
We conducted experiments to evaluate the effectiveness and robustness of our proposed hierarchical control framework for multi-step humanoid manipulation tasks on a real-world robot platform.
IV-A Experimental Setup
Experiments were conducted using the 29-degree-of-freedom (DoF) Unitree G1 humanoid robot. The robot was equipped with two externally mounted ELP RGB cameras, providing binocular egocentric visual inputs for the mid-level imitation learning (IL) policies. The robot was tested within a controlled environment consisting of two tables and objects arranged to simulate representative manipulation tasks common in household scenarios.
IV-B Evaluation Task
To validate the hierarchical framework, we defined a representative multi-step manipulation task involving:
  1. 1.
    Picking: The robot must grasp and lift the bag from the initial table.
  2. 2.
    Placing: The robot needs to accurately place the bag onto a second table.
IV-C Experimental Procedure
A total of 40 independent trials were conducted. The robot operated fully autonomously, executing plans generated by the high-level VLM planner, with skills dynamically selected and monitored for completion in real-time by the VLM execution monitor. Success was defined by the robot correctly picking up the bag, and successfully placing it on the designated target surface.
IV-D Experimental Results
Over 40 trials, the integrated hierarchical system successfully completed the entire manipulation sequence with an overall success rate of $72.5\%$. During each trial, we recorded failures at three distinct levels of the hierarchical framework. The observed failures were categorized as follows, in order of frequency:
  • Skill Policy Failure: Occurred when a mid-level imitation policy failed to execute the desired skill. Most frequently occurred during the grasping phase, primarily due to variations in object positioning beyond training data distributions.
  • VLM Execution Monitor Failure: Instances where the visual verification prematurely indicated task completion due to subtle positioning inaccuracies or ambiguous visual cues.
  • VLM Planner Failure: Occasional mis-grounding issues occurred, resulting in the generation of incorrect skill sequences.
Qualitatively, the robot exhibited stable and coherent manipulation behaviors across successful trials. Skills transitioned smoothly under guidance from the VLM execution monitor, demonstrating effective real-time adaptation and corrective feedback. Minor irregularities in the skills arose due to executing multiple consecutive 50-step action chunks, causing subtle positional resets between chunks. This suggests that introducing a smoothing or blending function between consecutive chunks could further enhance motion fluidity.
V Discussion
The experiments demonstrate that augmenting the canonical two-layer humanoid stack with a vision-language planning and monitoring layer enables reliable execution of multi-step non-prehensile manipulation routines in the physical world. Deployed on a 29-DoF Unitree G1, our proposed three-level architecture successfully completed the obstructed bag-transfer task in 72.5% of forty independent trials. The execution monitor’s one-second mean verification latency at 1Hz fits comfortably within the robot’s motion cycle, and coexists with high-frequency skill policies and RL controller, illustrating the large-model reasoning and perception need not preclude high-rate control.
Compared to prior humanoid manipulation systems, which often rely on open-loop sequences of predefined skills, our approach introduces continuous closed-loop reasoning. By integrating interpretable VLM-based planning with real-time visual verification, our framework dynamically selects appropriate skills, and explicitly determines when transitions between skills should occur.
Furthermore, unlike purely end-to-end vision-language-action (VLA) approaches, which depend heavily on large-scale, paired demonstration datasets and typically lack interpretability, our explicit planner-monitor design provides clear introspection into decision-making processes. This transparency facilitates targeted debugging, allowing developers to easily diagnose and address failures. Similarly, compared to strictly symbolic Task and Motion Planning (TAMP) frameworks, our system’s reliance on soft, natural-language-based skill descriptions—akin to a flexible version of PDDL—greatly simplifies the integration and extension of new behaviors without necessitating detailed domain-specific expertise.
V-A Failures and Limitations
A qualitative failure analysis reveals three dominant modes. First, skill-policy drift occurs when partial occlusions or out-of-distribution object poses cause the imitation policy to stall short of a grasp or placement. Second, planner mis-grounding emerges when the deliberative VLM hallucinates objects, or fails to sequence skills correctly. This typically involves adding additional irrelevant plan steps, or formulating redundant transition questions. Third, optimistic monitor judgments occasionally trigger a premature transition when subtle geometric conditions — such as a bag resting on the edge of a table — are mis-classified as success.
Several limitations temper the present findings. Our chosen evaluation task family, while non-trivial, is confined to table-top pick-place scenes and therefore under-samples the combinatorial diversity of domestic manipulation. The 1Hz verification rate, although adequate for deliberative motions, would be insufficient for fast skills, such as rapid recovery motions. This implicitly defines the timescale of skill abstraction.
V-B Future work
Our modular architecture provides clear pathways for future enhancements:
Error detection and adaptive re-planning:  A critical future enhancement includes adapting the VLM execution monitor system to explicitly detect execution failures or anomalous states. Through failure detection, we can trigger real-time adaptive re-planning. This promises to improve the robustness of the system.
Enhanced semantic and contextual conditioning:  Future work involves incorporating richer conditioning beyond visual and proprioceptive states from the high-level execution monitor. For example, replacing the ResNet [13] backbone with that of CLIP [25] would allow the VLM to have finer-grained semantic control of the skills via text-conditioning.
Expanding the skill library:  Finally, the current skill library is focused primarily on pick-and-place. An immediate next step involves expanding the robot’s skill repertoire. Examples include non-prehensile bi-manipulation and tool use. This broader skill-set would enable more complex, realistic long-horizon tasks, requiring the planning and monitoring components to handle increasingly diverse and dynamic environments. We’ve attempted to include the push skills that are triggered when an obstacle is in the way of the placing skill, as shown in Fig. 1, but we didn’t have time to finish a new set of experiments before submission.
VI Conclusion
In this work, we presented a hierarchical planning and control framework for multi-step humanoid manipulation tasks, combining robust RL-based whole-body tracking, imitation-learned skill policies, and a high-level vision-language planning and monitoring module. Evaluated on the Unitree G1 humanoid robot, our integrated system demonstrated successful autonomous execution of complex manipulation sequences, achieving a 72.5% success rate across real-world trials. Crucially, our approach offers inherent modularity, interpretability, and adaptability—enabling straightforward extension of the robot’s skill repertoire. This work highlights the effectiveness of integrating vision-language models with hierarchical control, paving the way toward increasingly capable autonomous humanoid systems operating in dynamic human environments.

### v3

@@
 This paper presents a hierarchical planning and control framework designed to achieve reliable multi-step humanoid manipulation.
-The proposed system comprises three layers: (1) a low-level RL-based controller responsible for tracking whole-body motion targets; (2) a mid-level set of skill policies trained via imitation learning that produce motion targets for different steps of a task; and (3) a high-level vision-language planning module that determines which skills should be executed and also monitors their completion in real-time using pretrained vision-language models (VLMs).
+The proposed system comprises three layers: (1) a low-level controller trained via reinforcement learning responsible for tracking whole-body motion targets, (2) a mid-level set of skill policies trained via imitation learning that produce motion targets for different steps of a task, and (3) a high-level vision-language planning module that determines which skills should be executed and also monitors their completion in real time using pretrained vision-language models (VLMs).
 Experimental validation is performed on a Unitree G1 humanoid robot executing a non-prehensile pick-and-place task.
-Over 40 real-world trials, the hierarchical system achieved a 72.5% success rate in completing the full manipulation sequence.
+Over 40 real-world trials, the hierarchical system achieved a 73% success rate in completing the full manipulation sequence.
 These experiments confirm the feasibility of the proposed hierarchical system, highlighting the benefits of VLM-based skill planning and monitoring for multi-step manipulation scenarios.
-See https://vlp-humanoid.github.io/ for video demonstrations of the policy rollout.
+See https://vlp-humanoid.github.io/ for a video demonstration of the policy rollout.
 I Introduction
@@
 We empirically validate our proposed hierarchical system on the 29-DoF Unitree G1 humanoid robot, performing a representative multi-step pick-and-place task.
-In real-world testing, the integrated system successfully completed the full manipulation sequence in 72.5% of 40 trials.
+In real-world testing, the integrated system successfully completed the full manipulation sequence in 73% of 40 trials.
 Our hierarchical approach demonstrates effective coordination of manipulation skills and visual feedback, highlighting the benefit of structured skill decomposition and VLM-based execution monitoring for reliable, multi-step humanoid autonomy.
@@
 These demonstrations were used to train the Humanoid Imitation Transformer (HIT)—a transformer-based imitation learning policy inspired by the Action Chunking Transformer (ACT) [31].
-Using only 20–40 demonstrations per skill, HumanPlus reported 60–100
+Using only 20–40 demonstrations per skill, HumanPlus reported 60–100% success rates across diverse whole-body manipulation tasks.
 Across these studies, a common hierarchical structure appears, featuring a robust RL-based low-level policy for tracking motion references and a high-level module generating these references through teleoperation or imitation learning.
@@
 Figure 2: Overview of the proposed hierarchical framework for autonomous multi-step humanoid manipulation.
-The system consists of: (1) a high-level planning module using vision-language models (VLMs) to generate task-specific skill sequences and monitor their execution via visual feedback; (2) a mid-level motion-generation module composed of imitation-learned action expert policies, mapping egocentric visual inputs and proprioceptive robot states into target joint configurations; and (3) a low-level reinforcement-learning-based tracking policy generating feasible joint-angle trajectories, executed by a high-frequency PD controller.
+The system consists of: (1) a high-level planning module using vision-language models (VLMs) to generate task-specific skill sequences and monitor their execution via visual feedback, (2) a mid-level motion-generation module composed of imitation-learned action expert policies, mapping egocentric visual inputs and proprioceptive robot states into target joint configurations, and (3) a low-level reinforcement-learning-based tracking policy generating feasible joint-angle trajectories, executed by a high-frequency PD controller.
 We propose a three-layer hierarchical control and planning framework for multi-step humanoid manipulation, illustrated in Fig.
@@
 III-A), (2) a mid-level motion-generation module composed of multiple imitation-learned action expert policies, each trained from teleoperated demonstrations (Sec.
-III-B);, and (3) a high-level vision-language-based planning and execution monitoring module (Sec.
+III-B), and (3) a high-level vision-language-based planning and execution monitoring module (Sec.
 III-C).
@@
 For the IL policies, we adopt the Humanoid Imitation Transformer (HIT) [10], designed specifically for high-DoF humanoids with binocular vision.
-It is based on the Action Chunking Transformer (ACT) [31] model, which have demonstrated improved robotic control by predicting multi-step future actions in one forward pass.
+It is based on the Action Chunking Transformer (ACT) [31] model, which has demonstrated improved robotic control by predicting multi-step future actions in one forward pass.
 At inference, HIT outputs action chunks of 50 joint-angle targets at 25 Hz.
@@
 containing synchronized state-action pairs recorded at 25 Hz.
-We collect data for the representative task rearrangement involving picking a bag from a table and and placing it onto another.
-Only successful sequences (30 in total) are retained, segmented into separate skill-specific datasets, i.e.
-picking and placing, for training two policies.
-IL policy training employs the demonstration data with hyperparameters specified in Table V.
-After training, we evaluate independently every skill policy, evaluation over 30 autonomous trials showed success rates of $90\%$ for picking and $83.3\%$ for placing.
+We collect data for the representative task rearrangement involving picking a bag from a table and placing it onto another.
+Only successful sequences (30 in total) are retained, segmented into separate skill-specific datasets, i.e., picking and placing, for training two policies.
+IL policy training employs the demonstration data with hyperparameters specified in Table VI.
+After training, we evaluated independently each skill policy over 30 autonomous trials, which showed success rates of $90\%$ for picking and $83\%$ for placing.
 III-C Multi-Step Skill Planning and Execution Monitoring
@@
 We represent each skill available to the robot using structured, human-readable descriptions that include the ‘preconditions’ and ‘effects’ fields of an operator definition in Planning Domain Definition Language (PDDL) [1].
-Each skill description includes a one-sentence summary, explicit natural-language preconditions, effects, and example verification queries, as illustrated for the Pick skill in Figure 3.
+Each skill description includes a one-sentence summary, explicit natural-language preconditions, effects, and example verification queries, as illustrated for the Pick skill in Fig.
+3.
 Unlike rigid symbolic representations, this PDDL-like form balances logical rigor and flexibility, enabling the planner to perform common-sense reasoning without exhaustive state enumeration.
@@
 III-D Full Hierarchical System
-The complete integrated system implements the hierarchical pipeline presented in Figure 2.
-The pipeline comprises: (1) a high-level VLM Planner (GPT-4o) that generates a sequence of skills from initial visual inputs and natural-language task instructions; (2) a real-time VLM Skill Monitor (Gemini-2.0-Flash-Lite) that verifies skill execution status based on visual feedback; (3) a Skill Selector that dynamically chooses the current skill to execute based on the planned sequence and monitor feedback; (4) a mid-level Imitation Learning Module that activates the corresponding imitation-learned policy according to the selected skill, producing joint-motion targets from egocentric images and proprioceptive data; and (5) a low-level Tracking Policy combined with a high-frequency PD Controller, converting these joint targets into executable motor torque commands.
+The complete integrated system implements the hierarchical pipeline presented in Fig.
+2.
+The pipeline comprises: (1) a high-level VLM planner (GPT-4o) that generates a sequence of skills from initial visual inputs and natural-language task instructions, (2) a real-time VLM skill monitor (Gemini-2.0-Flash-Lite) that verifies skill execution status based on visual feedback, (3) a skill selector that dynamically chooses the current skill to execute based on the planned sequence and monitor feedback, (4) a mid-level imitation learning Module that activates the corresponding imitation-learned policy according to the selected skill, producing joint-motion targets from egocentric images and proprioceptive data, and (5) a low-level tracking policy combined with a high-frequency PD controller, converting these joint targets into executable motor torque commands.
 A key advantage of this hierarchical design is its inherent modularity: individual IL skills are trained independently and added directly to the skill library, enabling effortless expansion of the robot’s behavioral repertoire.
@@
 The robot operated fully autonomously, executing plans generated by the high-level VLM planner, with skills dynamically selected and monitored for completion in real-time by the VLM execution monitor.
-Success was defined by the robot correctly picking up the bag, and successfully placing it on the designated target surface.
+Success was defined by the robot correctly picking up the bag and successfully placing it on the designated target surface.
 IV-D Experimental Results
-Over 40 trials, the integrated hierarchical system successfully completed the entire manipulation sequence with an overall success rate of $72.5\%$.
+TABLE I: Evaluation of individual policies and the full system.
+Pick Place Pick-Place
+Number of trials 30 30 40
+Number of successes 27 25 29
+Success rate 90% 83% 73%
+Over 40 trials, the integrated hierarchical system successfully completed the entire manipulation sequence with an overall success rate of $73\%$, as summarized in Table I.
 During each trial, we recorded failures at three distinct levels of the hierarchical framework.
@@
 The experiments demonstrate that augmenting the canonical two-layer humanoid stack with a vision-language planning and monitoring layer enables reliable execution of multi-step non-prehensile manipulation routines in the physical world.
-Deployed on a 29-DoF Unitree G1, our proposed three-level architecture successfully completed the obstructed bag-transfer task in 72.5% of forty independent trials.
-The execution monitor’s one-second mean verification latency at 1Hz fits comfortably within the robot’s motion cycle, and coexists with high-frequency skill policies and RL controller, illustrating the large-model reasoning and perception need not preclude high-rate control.
+Deployed on a 29-DoF Unitree G1, our proposed three-level architecture successfully completed the obstructed bag-transfer task in 73% of forty independent trials.
+The execution monitor’s one-second mean verification latency at 1Hz fits comfortably within the robot’s motion cycle, and coexists with high-frequency skill policies and RL controller, illustrating that the large-model reasoning and perception need not preclude high-rate control.
 Compared to prior humanoid manipulation systems, which often rely on open-loop sequences of predefined skills, our approach introduces continuous closed-loop reasoning.
-By integrating interpretable VLM-based planning with real-time visual verification, our framework dynamically selects appropriate skills, and explicitly determines when transitions between skills should occur.
+By integrating interpretable VLM-based planning with real-time visual verification, our framework dynamically selects appropriate skills and explicitly determines when transitions between skills should occur.
 Furthermore, unlike purely end-to-end vision-language-action (VLA) approaches, which depend heavily on large-scale, paired demonstration datasets and typically lack interpretability, our explicit planner-monitor design provides clear introspection into decision-making processes.
@@
 First, skill-policy drift occurs when partial occlusions or out-of-distribution object poses cause the imitation policy to stall short of a grasp or placement.
-Second, planner mis-grounding emerges when the deliberative VLM hallucinates objects, or fails to sequence skills correctly.
-This typically involves adding additional irrelevant plan steps, or formulating redundant transition questions.
-Third, optimistic monitor judgments occasionally trigger a premature transition when subtle geometric conditions — such as a bag resting on the edge of a table — are mis-classified as success.
+Second, planner misgrounding emerges when the deliberative VLM hallucinates objects or fails to sequence skills correctly.
+This typically involves adding additional irrelevant plan steps or formulating redundant transition questions.
+Third, optimistic monitor judgments occasionally trigger a premature transition when subtle geometric conditions—such as a bag resting on the edge of a table—are misclassified as success.
 Several limitations temper the present findings.
-Our chosen evaluation task family, while non-trivial, is confined to table-top pick-place scenes and therefore under-samples the combinatorial diversity of domestic manipulation.
+Our chosen evaluation task family, while non-trivial, is confined to table-top pick-and-place scenes and therefore under-samples the combinatorial diversity of domestic manipulation.
 The 1Hz verification rate, although adequate for deliberative motions, would be insufficient for fast skills, such as rapid recovery motions.
@@
 Our modular architecture provides clear pathways for future enhancements:
-Error detection and adaptive re-planning:  A critical future enhancement includes adapting the VLM execution monitor system to explicitly detect execution failures or anomalous states.
-Through failure detection, we can trigger real-time adaptive re-planning.
+Error detection and adaptive re-planning:  A critical future enhancement includes adapting the VLM execution monitor system to explicitly detect execution failures or anomalous states [23, 26].
+Through failure detection, we can trigger real-time adaptive replanning.
 This promises to improve the robustness of the system.
 Enhanced semantic and contextual conditioning:  Future work involves incorporating richer conditioning beyond visual and proprioceptive states from the high-level execution monitor.
-For example, replacing the ResNet [13] backbone with that of CLIP [25] would allow the VLM to have finer-grained semantic control of the skills via text-conditioning.
+For example, replacing the ResNet backbone with that of CLIP [24] would allow the VLM to have finer-grained semantic control of the skills via text-conditioning.
 Expanding the skill library:  Finally, the current skill library is focused primarily on pick-and-place.
@@
 Examples include non-prehensile bi-manipulation and tool use.
-This broader skill-set would enable more complex, realistic long-horizon tasks, requiring the planning and monitoring components to handle increasingly diverse and dynamic environments.
-We’ve attempted to include the push skills that are triggered when an obstacle is in the way of the placing skill, as shown in Fig.
-1, but we didn’t have time to finish a new set of experiments before submission.
+This broader skill set would enable more complex, realistic long-horizon tasks, requiring the planning and monitoring components to handle increasingly diverse and dynamic environments.
 VI Conclusion
 In this work, we presented a hierarchical planning and control framework for multi-step humanoid manipulation tasks, combining robust RL-based whole-body tracking, imitation-learned skill policies, and a high-level vision-language planning and monitoring module.
-Evaluated on the Unitree G1 humanoid robot, our integrated system demonstrated successful autonomous execution of complex manipulation sequences, achieving a 72.5% success rate across real-world trials.
+Evaluated on the Unitree G1 humanoid robot, our integrated system demonstrated successful autonomous execution of complex manipulation sequences, achieving a 73% success rate across real-world trials.
 Crucially, our approach offers inherent modularity, interpretability, and adaptability—enabling straightforward extension of the robot’s skill repertoire.